OpenStudy (anonymous):

–3(5 – 15n) = 15(3n – 2)

OpenStudy (compassionate):

Hey, So the first step is distributing your -3 and 15. Can you do that for me?

OpenStudy (anonymous):

i got 0=-15 ... is that no solution ?

OpenStudy (anonymous):

-3(5 - 15n) = 15(3n - 2) -15 + 45n = 45n - 30 -15 = -30 No solution.